# TESSERA: Precomputed FAIR Global Pixel Embeddings for Earth Representation and Analysis

### v5

@@ A.7 Downstream Task: Above-ground biomass estimation
For the AGB regression task, we used the ground truth data from the BioMassters dataset [72]. This dataset covers nearly 11,500 forest patches in Finland, with each patch representing a $\times2,560$  m area at 10 meter pixel resolution. Above-Ground Biomass (AGB) values were determined using open forest airborne LiDAR data and an extensive network of field plots as reference measurements by the Finnish Forest Centre. Patch coordinates and collection years–provided by the authors of Ref. [72] upon request–were used to retrieve satellite imagery to generate TESSERA and Google AlphaEarth Satellite Embeddings (GSE) [39].
@@
To obtain AGB estimates, we trained a convolutional neural network based on the UNet architecture [76], using TESSERA or GSE embeddings  [39] as input. The model received $\times256$$\times$128 and $\times256$$\times$64 input patches for TESSERA and GSE, respectively, where the dimension denotes the number of channels.
@@
Following the evaluation protocol in Ref. [72], RMSE was computed on a test set of 2773 patches and reported as the average RMSE per patch. For limited label testing, a subset of labels was randomly chosen from the remaining data and randomly divided into training and validation sets (80:20). The UNet model was trained on the embeddings for 80 epochs, with the model achieving the lowest loss on the validation set selected as the final model. Figure  5 shows the average RMSE calculated over five random splits, with bars indicating the minimum–maximum range when it exceeds the size of the dot. Training was performed with Adam Optimizer (learning rate = 1e-4, batch size = 4). Task-specific training on the entire dataset (80% training / 20% validation) using a single NVIDIA H200 GPU and 8 CPU cores took approximately 5 hours for TESSERA embeddings and 4 hours for the GSE embeddings.
@@
In addition to GSE, we compared the performance of TESSERA to two other baselines: the winner of the BioMassters competition [72] and the SpectralGPT foundation model [18] combined with a UPerNet [78].
@@
A.8 Downstream Task: Stocking indices in Voluntary Carbon Markets
Relative height modelling
Participants in the voluntary carbon market can use a remotely sensed “stocking index” to quantify project additionality, that is, the degree to which the carbon sequestration in the project exceeds what would have occurred in the natural course of events. The chosen stocking index must be correlated with the above-ground biomass (AGB) of the observed objects, and the strength of the correlation must be demonstrated using a suitable in situ reference dataset.
To demonstrate the superior quality of the TESSERA embeddings compared to existing state-of-the-art canopy height or AGB products, a simple stocking index was derived from GEDI relative heights and TESSERA embeddings. This permits comparing state-of-the-art models often used in the voluntary carbon market with a TESSERA-derived stocking index.
To model a suitable stocking index from the 128-dimensional TESSERA embeddings, we extracted 110,000 relative heights RH95 and RH10 from the GEDI Level 4A product [53] from 2022 from the land area within the 22MCT Sentinel-2 tile in Para Sate, Brazil. From the 10 m pixel falling at the centre of each full-waveform footprint, the 128 TESSERA embeddings were extracted as predictor variables for regression analysis, with the relative height differences (dRH = RH95-RH10) used as the target variables. A random forest (RF) regression model with 500 trees and 42 variables per split was used to model this stocking index. The TerraClass Amazonia 2022 land cover map [117] was used to distinguish between forest and non-forest. To account for implausible GEDI recordings often found in areas without significant tree cover (e.g., water surfaces or agricultural fields), the dRH values of these classes were set to zero before model training.
Stocking index baselines
For inference and assessment, the trained RF model was applied to eight Sentinel-2 tiles in the same region of Para State (MGRS tiles 22MCT, 22MCU, 22MDT, 22MDU, 22MGB, 22MGC, 22MHB, 22MHC). Within this area, a total of 38 agroforestry plots are located for which in situ measurements are available under a DTA. For comparison, five recent and widely used canopy height and AGB products were chosen and compared against the same set of 38 in situ AGB measurements: ETH CHM (10 m) [58], CTrees canopy tree height (5 m) [118], ESA-CCI AGB (100 m) [119], Meta CHM (1 m)  [60], and GTDX AGB (25 m) [120]. Using the same TerraClass Amazonia 2022 shapefiles that define the agroforestry systems, the pixels of the respective products were extracted and averaged for the chronosequence plots (n=38).
To ensure a fair comparison between canopy height, AGB and relative height products, the extracted averages were linearly rescaled to best correlate with in situ (AGB) data. After rescaling, all products are in units of t/ha and have an intercept of zero and a slope of one with respect to the in situ data, making their RMSE (and $R^{2}$) directly comparable. The rescaling is as follows:
  1. 1.
    TESSERA: difference in relative heights of GEDI estimates of RH90 and RH10 converted to AGB using 15.502 $\times$ + 160.5
  2. 2.
[…]
  3. 3.
    CTrees canopy height converted using 1.806 $\times$ + 44.9
  4. 4.
    ESA AGB converted using 0.407 $\times$ + 34.0
  5. 5.
    Meta canopy height converted using 3.723 $\times$ + 46.6
  6. 6.
    GEDI-Tandem-X AGB prediction converted using 0.3 $\times$ + 54.3
A.9 Interactive Habitat Mapping Tool
@@
The core of the tool is the interactive labelling workflow. A user can define custom habitat classes and provide training data by clicking on the map to place labelled points. These points and their corresponding 128-dimensional embedding vectors are used to train a k-Nearest Neighbors (kNN) classifier. The trained model then predicts a class for every pixel in the region of interest, producing a map layer that overlays a coregistered aerial view. This process allows for iterative refinement, where a user can add more points to correct misclassifications and retrain the model for immediate feedback, demonstrating an accessible method for quickly creating habitat maps without requiring deep machine learning expertise. We also provide a batch processing tool for larger regions of interest. The tools and their open source code are freely available in our github repository [37].
Table 5: TESSERA provides unprecedented ease of use, scale, and accuracy. Unlike other remote sensing foundation models, TESSERA provides analysis-ready and gap-free outputs (R), is open (O), and provides global (G) annual (A) coverage at 10 m resolution (10 m). It uses spectral-temporal features (ST) at pixel level (P), is invariant to clouds (CI), provides a compressed representation (CR), is oriented to land features (LO), and includes both multispectral and radar information (MS). Codes: $\checkmark$= Yes, $\times$= No, ? = Unknown, p= Potentially true, .. = N/A.
@@
TESSERA (2025) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
GeoKR (2021) [121] $\times$ ? $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$
CVPRW (2021) [122] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ .. $\checkmark$ $\times$
GASSL (2021) [23] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ .. $\checkmark$ $\times$
SeCo (2021) [123] $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ ? $\checkmark$ $\times$
MOSAIKS (2021) [124] $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ ?
DINO-MM (2022) [125] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ .. $\checkmark$ $\checkmark$
SatMAE (2022) [16] $\times$ .. $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$ p .. p $\times$
RS-BYOL (2022) [126] $\times$ .. .. $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ .. $\times$ $\checkmark$
GeCo (2022) [127] $\times$ .. .. $\times$ .. $\times$ $\times$ $\times$ .. p $\times$
RingMo (2022) [128] $\times$ .. .. $\times$ $\checkmark$ $\times$ $\times$ $\times$ ? $\checkmark$ $\times$
RVSA (2022) [129] $\times$ .. .. $\times$ .. $\times$ $\times$ $\times$ .. $\checkmark$ $\times$
RSP (2022) [14] $\times$ .. .. $\times$ p $\times$ $\times$ p $\checkmark$ $\checkmark$ $\times$
MATTER (2022) [130] $\times$ .. .. $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$
CSPT (2022) [131] $\times$ .. .. $\times$ ? $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$
CVPRW (2022) [132] $\times$ .. .. $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$
BFM (2023) [133] $\times$ .. .. $\times$ ? $\times$ $\times$ $\times$ ? $\checkmark$ $\times$
TOV (2023) [134] ? .. $\times$ $\times$ $\checkmark$ $\times$ ? $\times$ ? $\checkmark$ $\times$
CMID (2023) [135] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ ? $\checkmark$ $\times$
RingMo-Sense (2023) [136] $\times$ .. $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ ? $\checkmark$ $\checkmark$
lal-SimCLR (2023) [137] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$
CACo (2023) [138] $\times$ .. $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$
SatLas (2023) [15] $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$
GFM (2023) [139] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
Scale-MAE (2023) [140] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
DINO-MC (2023) [141] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
CROMA (2023) [142] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$
Cross-Scale MAE (2023) [143] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
DeCUR (2023) [144] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$
Presto (2023) [19] $\times$ .. $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
CtxMIM (2023) [145] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
FG-MAE (2023) [146] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
Prithvi (2023) [17] $\times$ .. $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$
RingMo-Lite (2023) [147] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
IGARSS (2023) [148] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
EarthPTi (2023) [149] $\times$ .. $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$
USat (2023) [150] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
AIEarth (2023) [151] p ? $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
Clay (2023) [152] $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
Hydro (2023) [153] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
U-Barn (2023) [154] $\times$ .. $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$
GeRSP (2023) [155] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
SwiMDiff (2023) [156] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
OFA-Net (2023) [157] $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$
SML-FR (2023) [158] p p $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
Spectral-GP (2024) [18] $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
S2MAE (2024) [159] $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
SatMAE++ (2024) [160] $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
msGFM (2024) [161] $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
SkySense (2024) [24] $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$
MTP (2024) [162] $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$
DOFA (2024) [163] $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$
MMEarth (2024) [164] $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$
LeMeViT (2024) [165] $\times$ .. $\times$ $\times$ ? $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\times$
SoftCon (2024) [166] $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$
RS-DFM (2024) [167] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$
A2MAE (2024) [168] $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$
HyperSIGM (2024) [169] $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$
SelectiveMAE (2024) [170] $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$
OmniSat (2024) [171] $\times$ .. $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ ? $\checkmark$ $\checkmark$ $\checkmark$
MM-VSE (2024) [172] $\times$ .. $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ ? $\checkmark$ $\checkmark$ $\checkmark$
MA3E (2024) [173] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$
SpectralEarth (2024) [174] $\times$ .. $\checkmark$ $\checkmark$ $\times$ ? $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$
SenPa-MAE (2024) [175] $\times$ .. $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$
RingMo-Ae (2024) [176] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$
SAR-JEPA (2024) [177] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$
PIS (2024) [178] $\times$ ? $\checkmark$ $\checkmark$ $\checkmark$ .. .. .. .. .. ..
OReole-FM (2024) [179] $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$
PIEVIT (2024) [180] $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$
SatVisionTOA (2024) [181] $\times$ .. $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$
RS-vHeat (2024) [182] $\times$ .. ? ? ? ? $\checkmark$ ? $\checkmark$ ? $\checkmark$
Prithvi-EO-2.0 (2024) [110] $\times$ .. $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$
AnySat (2024) [183] $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$
WildSAT (2024) [184] $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$
SeaMo (2024) [185] $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$
FoMo (2025) [186] $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$
SatMamba (2025) [187] $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
Galileo (2025) [188] p p $\checkmark$ ? $\checkmark$ $\times$ ? ? $\checkmark$ $\checkmark$ $\checkmark$
SatDiFuser (2025)[189] $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
RoMA (2025) [190] $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
Panopticon (2025) [191] $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$
FedSense (2025) [192] $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
Copernicus-FM (2025) [193] $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$
HyperFree (2025) [194] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$
DynamicVis (2025) [195] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$
FlexiMo (2025) [196] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$
Google Satellite (2025)[39] $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
SkySense++ (2025) [197] $\times$ $\checkmark$ $\checkmark$ p p $\checkmark$ $\times$ p $\checkmark$ $\checkmark$ $\checkmark$
RingMoE (2025) [198] $\times$ .. $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$

### v6

@@ A.9 Downstream Task: Above-ground biomass estimation
For the AGB regression task, we used the ground truth data from the BioMassters dataset [63]. This dataset covers nearly 11,500 forest patches in Finland, with each patch representing a $\times2,560$  m area at 10 meter pixel resolution. Above-Ground Biomass (AGB) values were determined using open forest airborne LiDAR data and an extensive network of field plots as reference measurements by the Finnish Forest Centre. Patch coordinates and collection years–provided by the authors of Ref. [63] upon request–were used to retrieve satellite imagery to generate TESSERA and Google AlphaEarth Satellite Embeddings (AlphaEarth) [11].
@@
To obtain AGB estimates, we trained a convolutional neural network based on the UNet architecture [73], using TESSERA or AlphaEarth embeddings  [11] as input. The model received $\times256$$\times$128 and $\times256$$\times$64 input patches for TESSERA and AlphaEarth, respectively, where the dimension denotes the number of channels.
@@
Following the evaluation protocol in Ref. [63], RMSE was computed on a test set of 2773 patches and reported as the average RMSE per patch. For limited label testing, a subset of labels was randomly chosen from the remaining data and randomly divided into training and validation sets (80:20). The UNet model was trained on the embeddings for 80 epochs, with the model achieving the lowest loss on the validation set selected as the final model. Figure  5 shows the average RMSE calculated over five random splits, with bars indicating the minimum–maximum range when it exceeds the size of the dot. Training was performed with Adam Optimizer (learning rate = 1e-4, batch size = 4). Task-specific training on the entire dataset (80% training / 20% validation) using a single NVIDIA H200 GPU and 8 CPU cores took approximately 5 hours for TESSERA embeddings and 4 hours for the AlphaEarth embeddings.
@@
In addition to AlphaEarth, we compared the performance of TESSERA to two other baselines: the winner of the BioMassters competition [63] and the SpectralGPT foundation model [34] combined with a UPerNet [108].
@@
A.10 Interactive Habitat Mapping Tool
@@
The core of the tool is the interactive labelling workflow. A user can define custom habitat classes and provide training data by clicking on the map to place labelled points. These points and their corresponding 128-dimensional embedding vectors are used to train a k-Nearest Neighbors (kNN) classifier. The trained model then predicts a class for every pixel in the region of interest, producing a map layer that overlays a coregistered aerial view. This process allows for iterative refinement, where a user can add more points to correct misclassifications and retrain the model for immediate feedback, demonstrating an accessible method for quickly creating habitat maps without requiring deep machine learning expertise. We also provide a batch processing tool for larger regions of interest. The tools and their open source code are freely available in the github repository.
A.11 Pseudo Code
Algorithm 1 Mixup Consistency Regularization for Multimodal Barlow Twins
1:def MixupBT(s2_aug1, s2_aug2, s1_aug1, s1_aug2, model, barlow_lambda, lambd_mix):
2:# s2_* and s1_* are batches of augmented inputs
3:# model returns projection z and representation repr
4:# standard Barlow Twins branch
5:z1, repr1 = model(s2_aug1, s1_aug1)
6:z2, repr2 = model(s2_aug2, s1_aug2)
7:loss_bt, bar_bt, off_bt = BarlowTwins(z1, z2)
8:# mixup branch
9:B = s2_aug1.shape[0]
10:idx = randperm(B)
11:alpha = Beta(beta_alpha, beta_beta).sample()
12:# mix S2 and S1 views
13:y_m_s2 = alpha * s2_aug1 + (1 - alpha) * s2_aug2[idx]
14:y_m_s1 = alpha * s1_aug1 + (1 - alpha) * s1_aug2[idx]
15:z_m, repr_m = model(y_m_s2, y_m_s1)
16:# cross-correlation of mixed and original views
17:cc_m_a = Corr(z_m, z1)
18:cc_m_b = Corr(z_m, z2)
19:cc_11 = Corr(z1, z1)
20:cc_2idx1 = Corr(z2[idx], z1)
21:cc_12 = Corr(z1, z2)
22:cc_2idx2 = Corr(z2[idx], z2)
23:# mixup targets and consistency loss
24:cc_m_a_gt = alpha * cc_11 + (1 - alpha) * cc_2idx1
25:cc_m_b_gt = alpha * cc_12 + (1 - alpha) * cc_2idx2
26:diff_a = (cc_m_a - cc_m_a_gt).pow(2).sum()
27:diff_b = (cc_m_b - cc_m_b_gt).pow(2).sum()
28:loss_mix = lambd_mix * barlow_lambda * (diff_a + diff_b)
29:# final SSL loss with mixup consistency
30:loss = loss_bt + loss_mix
31:return loss
Algorithm 2 Global Shuffling and Multimodal Data Construction
1:def GlobalShuffling(tile_roots, batch_size, T):
2:# tile_roots: list of paths to MGRS tile directories
3:# batch_size: number of tiles to process in memory simultaneously
4:# T: number of time steps for sparse sampling
5:tiles = shuffle(tile_roots) # Randomize tile processing order
6:num_batches = len(tiles) // batch_size
7:for b in range(num_batches):
8:    current_tiles = tiles[b*batch_size : (b+1)*batch_size]
9:    pixel_buffer_aug1 = []
10:    pixel_buffer_aug2 = []
11:    # Parallel processing of tiles to extract valid pixels
12:    parfor tile in current_tiles:
13:      s2, s1, masks = load_data(tile)
14:      valid_locs = filter_valid_pixels(masks, s1, s2)
15:      if len(valid_locs) == 0: continue
16:      # Sparse random temporal sampling per pixel
17:      for (y, x) in valid_locs:
18:        idx = get_valid_time_indices(masks, y, x)
19:        v1 = Sample(s2, s1, idx, T) # View 1
20:        v2 = Sample(s2, s1, idx, T) # View 2
21:        pixel_buffer_aug1.append(v1)
22:        pixel_buffer_aug2.append(v2)
23:    # Aggregate pixels from spatially disjoint tiles
24:    stack_aug1 = Stack(pixel_buffer_aug1) # Shape: (N_total, T, C)
25:    stack_aug2 = Stack(pixel_buffer_aug2)
26:    # Global Shuffling: Decorrelate spatial neighborhoods
27:    N_total = stack_aug1.shape[0]
28:    perm_idx = randperm(N_total)
29:    shuffled_aug1 = stack_aug1[perm_idx]
30:    shuffled_aug2 = stack_aug2[perm_idx]
31:    # Save as chunks for pre-training dataloader
32:    save_chunks(shuffled_aug1, shuffled_aug2, out_dir)
Table 9: TESSERA provides unprecedented ease of use, scale, and accuracy. Unlike other remote sensing foundation models, TESSERA provides analysis-ready and gap-free outputs (R), is open (O), and provides global (G) annual (A) coverage at 10 m resolution (10 m). It uses spectral-temporal features (ST) at pixel level (P), is invariant to clouds (CI), provides a compressed representation (CR), is oriented to land features (LO), and includes both multispectral and radar information (MS). Codes: Y= Yes, N= No, ? = Unknown, P= Potentially true, .. = N/A.
@@ A.11 Pseudo Code
TESSERA (2025) Y Y Y Y Y Y Y Y Y Y Y
GeoKR (2021) [44] N ? N N Y N Y N N Y N
CVPRW (2021) [80] N .. N N Y N N N .. Y N
GASSL (2021) [6] N .. N N Y N Y N .. Y N
SeCo (2021) [59] N N N N Y N N N ? Y N
MOSAIKS (2021) [72] N Y Y N N N N N Y Y ?
DINO-MM (2022) [98] N .. N N Y N N N .. Y Y
SatMAE (2022) [15] N .. N N Y Y N P .. P N
RS-BYOL (2022) [39] N .. .. N Y N Y Y .. N Y
GeCo (2022) [45] N .. .. N .. N N N .. P N
RingMo (2022) [81] N .. .. N Y N N N ? Y N
RVSA (2022) [93] N .. .. N .. N N N .. Y N
RSP (2022) [92] N .. .. N P N N P Y Y N
MATTER (2022) [2] N .. .. N Y Y N N Y Y N
CSPT (2022) [117] N .. .. N ? N N N Y Y N
CVPRW (2022) [75] N .. .. N Y N N N Y Y Y
BFM (2023) [12] N .. .. N ? N N N ? Y N
TOV (2023) [85] ? .. N N Y N ? N ? Y N
CMID (2023) [62] N .. N N Y N N N ? Y N
RingMo-Sense (2023) [112] N .. N N Y Y N N ? Y Y
lal-SimCLR (2023) [67] N .. N N Y N N N Y Y Y
CACo (2023) [57] N .. N N Y Y N N Y Y N
SatLas (2023) [8] N .. Y N Y Y N N Y Y N
GFM (2023) [61] N .. N N Y N N N N Y N
Scale-MAE (2023) [71] N .. N N Y N N N N Y N
DINO-MC (2023) [105] N .. N N Y N N N N Y N
CROMA (2023) [29] N .. N N Y N N N N Y Y
Cross-Scale MAE (2023) [84] N .. N N Y N N N N Y N
DeCUR (2023) [101] N .. N N Y N N Y N Y Y
Presto (2023) [88] N .. Y Y Y Y Y Y Y Y Y
CtxMIM (2023) [116] N .. N N Y N N N N Y N
FG-MAE (2023) [99] N .. N N Y N N N N Y N
Prithvi (2023) [76] N .. N N N Y N Y Y Y N
RingMo-Lite (2023) [100] N .. N N Y N N N N Y N
IGARSS (2023) [27] N .. N N Y N N N N Y N
EarthPTi (2023) [78] N .. N N Y Y Y Y Y Y N
USat (2023) [38] N .. N N Y N N N N Y N
AIEarth (2023) [111] P ? Y N Y N N N N Y N
Clay (2023) [14] N .. Y N Y N N N N Y N
Hydro (2023) [17] N .. N N Y N N N N N N
U-Barn (2023) [24] N .. N Y Y Y N Y N Y N
GeRSP (2023) [36] N .. N N Y N N N N Y N
SwiMDiff (2023) [86] N .. N N Y N N N N Y N
OFA-Net (2023) [110] N .. Y N Y N N N N Y Y
SML-FR (2023) [21] P P Y N Y N N N N N N
Spectral-GP (2024) [34] N N Y N Y N N N N Y N
S2MAE (2024) [47] N N Y N Y N N N N Y N
SatMAE++ (2024) [65] N N Y N Y N N N N Y N
msGFM (2024) [32] N N N N Y N N N N Y N
SkySense (2024) [31] N N N N Y Y N N N Y Y
MTP (2024) [94] N .. Y N Y N Y N N Y N
DOFA (2024) [109] N .. Y N Y N N N Y Y Y
MMEarth (2024) [64] N .. Y N Y N N N Y Y Y
LeMeViT (2024) [41] N .. N N ? N N N Y N N
SoftCon (2024) [102] N .. Y N Y N Y N N Y N
RS-DFM (2024) [104] N .. N N Y N Y N N Y N
A2MAE (2024) [115] N .. Y N Y N N N Y Y N
HyperSIGM (2024) [95] N .. Y N Y Y Y N Y Y N
SelectiveMAE (2024) [96] N .. Y N Y N N N Y Y N
OmniSat (2024) [4] N .. N Y Y Y N ? Y Y Y
MM-VSE (2024) [70] N .. Y Y Y Y N ? Y Y Y
MA3E (2024) [51] N .. N N Y N N N Y Y N
SpectralEarth (2024) [50] N .. Y Y N ? Y N Y Y N
SenPa-MAE (2024) [68] N .. Y Y Y Y Y N Y Y N
RingMo-Ae (2024) [19] N .. N N Y N Y N Y Y N
SAR-JEPA (2024) [46] N .. N N Y N N Y N N N
PIS (2024) [3] N ? Y Y Y .. .. .. .. .. ..
OReole-FM (2024) [20] N .. Y N Y N Y N Y Y N
PIEVIT (2024) [55] N .. Y N Y N Y N Y Y N
SatVisionTOA (2024) [79] N .. Y Y N Y Y N Y Y N
RS-vHeat (2024) [35] N .. ? ? ? ? Y ? Y ? Y
Prithvi-EO-2.0 (2024) [82] N .. Y Y N Y N Y Y Y N
AnySat (2024) [5] N .. Y N Y N Y N Y Y Y
WildSAT (2024) [18] N N N N N N N N Y Y N
SeaMo (2024) [49] N N N N Y N N N Y Y N
FoMo (2025) [10] Y Y Y N Y N Y N Y Y Y
SatMamba (2025) [23] Y Y N N N N N N N Y N
Galileo (2025) [89] P P Y ? Y N ? ? Y Y Y
SatDiFuser (2025)[40] N N N N Y N N N N Y N
RoMA (2025) [97] N N N N Y N N N N Y N
Panopticon (2025) [91] N Y N N Y Y N N Y Y Y
FedSense (2025) [83] N N N N N N N N N Y N
Copernicus-FM (2025) [103] N .. Y N Y Y N N Y Y Y
HyperFree (2025) [43] N .. N N Y N N N Y Y N
DynamicVis (2025) [13] N .. N N Y N N N N Y Y
FlexiMo (2025) [48] N .. N N Y N N N N Y Y
Google Satellite (2025)[11] Y N Y Y Y Y N Y Y Y Y
SkySense++ (2025) [106] N Y Y P P Y N P Y Y Y
RingMoE (2025) [9] N .. N N Y Y N N Y Y Y